OpenStudy (anonymous):

Please Help! Consider the function. f (x) = 1/2x +1 a. Find the inverse of f(x) and name it g(x). Show and explain your work. b. Use function composition to show that f(x) and g(x) are inverses of each other. c. Draw the graphs of f(x) and g(x) on the same coordinate plane. Explain how your graph shows that the functions are inverses of each other.

OpenStudy (anonymous):

how would you solve \[\frac{1}{2}x+1=7\]? a) subtract 1 b) multiply by 2 that is what the inverse does \[g(x)=(x-1)\times 2=2x-2\]

OpenStudy (anonymous):

ok i kinda get it. What about question B? can you help me with that?